# Clustering MOOC Programming Solutions to Diversify Their Presentation to Students

### v1

@@ 6. Dataset
0 (0-10] (10-50] (50-90] (90-100] 100
Tasks 387 134 79 109 112 46
% 44.6 15.5 9.1 12.6 12.9 5.3
@@
7. Fine-tuning parameters
@@
For the fine-tuning, we took the 17 tasks as described in the previous section, and for each of them, obtained 5 different solutions for showing using either JPlag or Rhubarb with different parameters, respectively. For example, for JPlag, we compared four different values of the threshold, so for each of the 17 tasks, we obtained four different groups of five solutions and rated them. The rating was carried out on the 1-to-5 scale, with 1 indicating the least suitable sample in terms of the diversity, code quality, and usefulness and 5 indicating the most suitable sample in terms of the diversity, code quality, and usefulness. Fine-tuning of the parameters was carried out by the first two authors of the paper, who have several years of experience working with MOOCs and programming in Python, and who labeled all 17 tasks. The final score for each setting was averaged between all 17 tasks and both labelers.
@@
JPlag
Threshold 0.2 0.4 0.6 0.8
Average score 3.85 3.82 3.82 3.71
Rhubarb
Distance limit 15 30 45 60
Average score 3.15 3.06 3.06 3.24
Distance limit 75 90 105 120
Average score 3.09 3.03 2.82 2.74
@@ 8. Comparison
Table 3. An average score of the default platform approach, JPlag, and Rhubarb.
Platform JPlag Rhubarb
Average score 3.12 3.77 3.50
The averaged results are presented in Table 3. As evident from the obtained data, both JPlag and Rhubarb offer more diverse solutions than those currently displayed on the Hyperskill platform. This indicates the usefulness of clustering in this context and of developing a new approach for this task.
On the one hand, the evaluation has shown that JPlag works better than Rhubarb, while both of them work better than the default approach. On the other hand, the comparison itself was carried on just a small sample of the data, since JPlag fails for the majority. The Rhubarb tool handled 100% of the dataset and can produce clusters with better quality than the default platform’s solution for all tasks. Importantly, JPlag can not be extended to all cases, since greedy string tiling algorithms by their definition can not be used for short submissions, while the quality of Rhubarb can be potentially improved. For now, we designed a system that employs JPlag on the 5.3% of tasks it can fully process, and employs Rhubarb on the remaining tasks.
@@ 9. Conclusion and Future Work
This work introduces a new problem of selecting algorithmically diverse solutions in order to display them to students on MOOCs. To address it, we proposed a tool called Rhubarb (Artser et al., 2024b) that brings Python solutions to a standard form, calculating distance via the GumTree tool that takes into account code structure, and selecting final examples based on code quality. The part of Rhubarb responsible for code transformations can be reused as a library (Artser et al., 2024a).
Using the data provided to us by the Hyperskill MOOC platform (jet, [n. d.]), we compared the default platform approach of simply showing the last student submissions, the adapted existing plagiarism tool JPlag, and Rhubarb. In the manual labelling by experts, the default approach received the average score of 3.12 out of 5, JPlag — 3.77, and our tool — 3.50. The comparison was performed only on a small sample of the data, since JPlag cannot fully process the majority of our tasks, with 44.6% not being processed at all. Like other state-of-the-art approaches, JPlag uses a greedy string tiling algorithm, the main limitation of which is the minimum length of solutions that can be handled. For Python, this limitation is critical, because even complex Python tasks can have short solutions (e.g., solution (e) from Figure 1). Rhubarb looks promising, because it can handle 100% of the data, and works better than the default approach. In the end, we created a compound version that uses both tools together.
Our future work relates to both research and practice. From the research point of view, we could increase the quality of Rhubarb by improving the GumTree processing and trying different types of clustering, e.g., mean-shift algorithms for clustering (Wu and Yang, 2007). Also, given the exploratory nature of the presented work, it is imperative to conduct a more thorough evaluation of hyperparameters.
From the practical standpoint, it is critical to make sure that the system actually helps students, since in this work we only focused on the initial exploration of the novel problem. We plan to actually incorporate the developed system into the Hyperskill platform, since it outperforms the current version, and conduct studies into the effect of showing alternative solutions to students. In addition to validating our approach, this will allow to create a more general framework for exploring this problem, comparing different techniques, and collecting student feedback on other forms of diversifying their knowledge that might not be obvious.
We hope that this work can bring more researchers into this under-developed problem so that MOOC platforms can serve students even better.

### v2

@@ 6.1. Dataset
[…]
The absolute and relative number of tasks, for which JPlag can process different percentage of solutions. E.g, the column (10-50] indicates the number of tasks, for which 10-50% of all submissions were present in JPlag’s output. 0 (0-10] (10-50] (50-90] (90-100] 100 № of tasks 387 134 79 109 112 46 % of tasks 44.6 15.5 9.1 12.6 12.9 5.3
@@
6.2. Fine-tuning Parameters
@@
For the fine-tuning, we took the 17 tasks as described in the previous section, and for each of them, obtained 5 different solutions for showing using either JPlag or Rhubarb with different parameters, respectively. For example, for JPlag, we compared four different values of the threshold, so for each of the 17 tasks, we obtained four different groups of five solutions and rated them. The rating was carried out on the 1-to-5 scale, with 1 indicating the least suitable sample in terms of the diversity, code quality, and usefulness and 5 indicating the most suitable sample in terms of the diversity, code quality, and usefulness. A single score was given, reflecting an overall assessment of all three dimensions combined. The fine-tuning of the parameters was carried out by the first two authors of the paper, who have several years of experience working with MOOCs and programming in Python, and who labeled all 17 tasks. The final score for each setting was averaged between all 17 tasks and both labelers.
@@
[…]
The results of fine-tuning for JPlag threshold and Rhubarb distance limit. JPlag Rhubarb Threshold 0.2 0.4 0.6 0.8 Dist. limit 15 30 45 60 75 90 105 120 Avg. score 3.85 3.82 3.82 3.71 Avg. score 3.15 3.06 3.06 3.24 3.09 3.03 2.82 2.74
@@ 6.3. Comparison
The averaged results are: Platform – 3.12, JPlag – 3.77, Rhubarb – 3.50. As evident from the obtained data, both JPlag and Rhubarb offer more diverse solutions than those currently displayed on the Hyperskill platform. This indicates the usefulness of clustering in this context and of developing a new approach for this task.
On the one hand, the evaluation has shown that JPlag works better than Rhubarb, while both of them work better than the default approach. On the other hand, the comparison itself was carried on just a small sample of the data, since JPlag fails for the majority. The Rhubarb tool handled 100% of the dataset and can produce clusters with better quality than the default platform’s solution for all tasks. Importantly, JPlag can not be extended to all cases, since greedy string tiling algorithms by their definition can not be used for short submissions, while the quality of Rhubarb can be potentially improved, and it is already rather close to JPlag. For now, we designed a system that employs JPlag on the 5.3% of tasks it can fully process, and uses Rhubarb on the remaining tasks.
7. Discussion
With this work, we aimed to explore the novel task of selecting diverse solutions to programming tasks for showing to students, both for research purposes and for using in a real MOOC platform. While we only conducted preliminary fine-tuning and pilot evaluation, some general findings are worth discussing.
Comparison to the default approach. Firstly and most obviously, it can be seen that both clustering approaches outperform the default platform approach of showing the latest solutions. This indicates the clear room for improvement for the existing implementation and the potential of the proposed concept in general.
Comparison to JPlag. At the same time, among the clustering-based approaches, the existing tool JPlag demonstrated better performance than Rhubarb. While this does not make Rhubarb obsolete due to it being able to handle all the submissions, it does show the need to integrate other distance measuring and clustering techniques into our approach. In the future, different methods should be combined and evaluated for this task.
Overall scores. Finally, one could argue that all the obtained scores are average at best, raising more research questions about showing other students’ solutions in the first place. We argue that this does not decrease the value of the proposed problem and our work in particular. Firstly, other ways of teaching different ways of solving problems (e.g., pre-written teachers’ solutions) are not always feasible on MOOCs. Secondly, the system of showing other students’ solutions is already implemented in several major MOOCs and thus already affects thousands of students. What this does indicate is the necessity to study this aspect of MOOCs more deeply, in particular, by comparing the proposed approaches in real life, by running A/B experiments with real students on MOOCs.
We plan to do this in our future work, and we hope that this paper can serve as an important first step towards bringing the attention of other researchers to this aspect of MOOCs and its practical value.
@@ 8. Conclusion and Future Work
This work introduces a new problem of selecting algorithmically diverse solutions in order to display them to students on MOOCs. To address it, we proposed a tool called Rhubarb (Artser et al., 2024b) that brings Python solutions to a standard form, calculates edit distances between them via the GumTree tool, and selects final examples based on code quality. Using the data provided to us by the Hyperskill MOOC platform (jet, 2024), we compared the default platform approach of simply showing the last student submissions, the adapted existing plagiarism tool JPlag, and Rhubarb. In the manual labelling by experts, the default approach received the average score of 3.12 out of 5, JPlag — 3.77, and our tool — 3.50. The comparison was performed only on a small sample of the data, since JPlag cannot fully process the majority of our tasks, with 44.6% not being processed at all. In the end, we created a compound version that uses both tools together to handle 100% of the data with good quality of the output that is higher than the standard platform’s approach.
Future work. Our future work relates to both research and practice. From the research point of view, we could increase the quality of Rhubarb by improving the GumTree processing and trying different types of clustering, e.g., mean-shift algorithms for clustering (Wu and Yang, 2007). Also, it is imperative to conduct a more thorough evaluation of hyperparameters, and in general — carry out more expert labeling. From the practical standpoint, it is critical to make sure that the system actually helps students, since in this work we only focused on the initial exploration of the problem. We plan to actually incorporate the developed system into the studied MOOC platform, since it outperforms the current version, and conduct studies into the effect of showing alternative solutions to students. In addition to validating our approach, this will allow to create a more general framework for exploring this problem, comparing different techniques, and collecting student feedback on other forms of diversifying their knowledge that might not be obvious.